In 7 years Deac will be twice as old as Irina. Four years ago he was three times as old as Irina was then. How old are each of them now?
I can not figure out how to set up the equations! The book says the answers are: Deac is 37 and Irina is 15. I do not come up with those ages. I have been using: x = Deac, y = Irina x + 7 = 2y and x - 4 = 3y. Help! What am I doing wrong? Thank you - Janette

Answer by rapaljer(4671)

You are correct in your setup, except for one point--you forgot that Irina's age is changing as well as Deac's age.
x= Deac's age now
y= Irina's age now

x+7= 2(y+7)
x-4= 3(y-4)

I recommend that you solve for x in both of the above equations, and set them equal to each other in order to solve for y. I get y= 15, which seems correct.

R^2 at SCC